OpenStudy (anonymous):

Find A, B, and C in the quadratic function (y=ax^2+bx+c) with the points (0,0) (1, 130) (2, 240) (11, ?) (If possible explain how to get it also)

OpenStudy (science0229):

First, it passes through (0,0), right?

OpenStudy (science0229):

So, 0=a(0)^2+b(0)+c=c Therefore, c=0

OpenStudy (anonymous):

yeah but I need tp figure out what A and B are also

OpenStudy (science0229):

I know. We'll get there, soon enough

OpenStudy (shoba):

Now you have coordinates for three points: (0, 0), (1, 130), and 2, 240). Plug them into the quadratic equation ax² + bx + c = y to get a system of three equations: 0²a + 0b + c = 0 1²a + 1b + c = 130 2²a + 2b + c = 240 Solve the system to get a = -10 b = 140 c = 0 -10x² + 140x = y equation of the parabola. A1 11 , y = -10·11² + 140·11 = 330

OpenStudy (science0229):

Well, that's your answer.

OpenStudy (anonymous):

I mainly can't figure out how to solve the system :/ I have like 3 more questions like this, just with different numbers